OpenStudy (anonymous):

write the logarithmic equation in exponential form a) log3 81 = 4 b) log6 216 = 3

OpenStudy (anonymous):

\[\log_{a}(y)=x \] \[y=a ^{x}\] Go!

OpenStudy (anonymous):

Log3 81/81=4/81 Log 3=4/81 10^(log3)=10(4/10) 3=10(4/81) Like that?

OpenStudy (anonymous):

Nah, \[81=3^4\]

OpenStudy (anonymous):

ok, i think im trying to make this way to hard. Thanks